OpenStudy (anonymous):

What is the value of the function when x = 4? y = 3x + 7 A. 38 B. 19 C. 76

OpenStudy (anonymous):

y = 3(4) + 7 y = 12 + 7 y = 19

OpenStudy (anonymous):

typo lol

OpenStudy (anonymous):

@tburn25 thankkssss